OpenStudy (anonymous):

"June is two years older than Sarah, and Sarah is four years younger than Hal. In 4 years, Hal's age will be equal to the sum of the present ages of Sarah and June. Find the age of each." How do I solve this?

OpenStudy (amistre64):

Have you tried yelling at it?

OpenStudy (anonymous):

A few times. It's pretty hard headed.

OpenStudy (amistre64):

j = s+2 s = h-4 so far :)

OpenStudy (amistre64):

h+4 = j+s

OpenStudy (amistre64):

h = s+4 s+4+4 = s+2 + s+4-4

OpenStudy (amistre64):

s+8 = 2s +2 6 = s ; but lets dbl chk that lol

OpenStudy (amistre64):

j = 8; s = 6; h = 10 to start with in 4 years h = 14 14 = 8+6; got it :)

OpenStudy (anonymous):

The question is basically saying: June is 2 years younger than Hal==> J=H-2 Sara is 4 years younger than Hal==>S=H-4 In four years, Hal will be four years older==> H+4 which will be equal to the sum of Sara and June's current age Therefore, H+4=H-2+H-4 Solve for H, You get, 10 Plug in 10 for H in June and Sara's equation from the first year You get: Sara=H-4 = 10-4 = 6 June=H-2 = 10-2 =8